# Texture Map Optimization and Physically-Based Rendering

### v1

@@ 1 Introduction
To reduce such burdens, methods for generating diverse 3D assets have been extensively studied [39, 58, 12, 34, 40, 9, 18]. Notably, the recent progress in neural volumetric representations, e.g., NeRF [38] and generative diffusion models [45, 46] have advanced the development of text-driven 3D asset generation [29, 41, 27, 33], which leverages a cheaper guidance, i.e., text description. While these methods generate coarse geometries and textures, the qualities are still unsatisfactory. Moreover, to use the generated assets in real graphics engines, e.g., Blender or Unity, one must convert the implicit volumetric geometries and textures into the explicit mesh and compatible texture maps, which makes them impractical. Manual extraction of mesh surfaces and unwrapping of textures could be performed, but it still has limitations. The unwrapped texture maps inevitably have heterogeneous texture mappings, so we cannot easily transfer or edit them, which could be the bottleneck for generating diverse 3D assets.
@@
In experiments, we demonstrate that Paint-it produces high-quality texture maps for general 3D meshes, e.g., objects, humans, and animals (Fig. 1). Also, Paint-it synthesizes a remarkable quality of texture maps compared to competing methods. As a favorable byproduct, the synthesized PBR texture maps are compatible with the popular graphics engine and can be seamlessly integrated into relighting and material control pipelines (Fig. 2).
We summarize our main contributions as follows:
@@ 2 Related Work
Our task is related to text-driven 3D generation, focusing mainly on the task of text-driven texture map synthesis. We briefly review these lines of work.
@@ 3 Paint-it: Text-Driven PBR Texture Synthesis via Neural Re-parameterized Optimization
In this section, we introduce our text-driven realistic mesh texture synthesis pipeline, Paint-it. We briefly overview the Score-Distillation Sampling, the loss for our optimization (Sec. 3.1). Then, we introduce the goal of the Paint-it optimization (Sec. 3.2), and details of our deep convolutional texture neural re-parameterization, DC-PBR (Sec. 3.3). Finally, we explain our text-driven physically-based rendering texture map optimization (Sec. 3.4).
Figure 3: Paint-it: Overall pipeline. Given a 3D object mesh without a texture and a text describing the desired appearance of the mesh, Paint-it synthesizes realistic PBR texture maps via synthesis-through-optimization. We introduce DC-PBR, which parameterizes the PBR texture map into randomly initialized convolutional neural kernels. By performing texture mapping to texturize the given mesh, we differentiably rasterize the textured mesh and obtain multi-view images. Using a diffusion-guided loss, Paint-it optimizes the neural parameters of the DC-PBR rather than directly optimizing the pixel values of the texture map.
@@ 3.3 DC-PBR: Deep Convolutional PBR Texture Map Re-parameterization
We propose the deep convolutional re-parameterization of PBR texture maps, DC-PBR, $\mathcal{T}_{\theta}$. Instead of the pixel value parameterization of texture maps, using DC-PBR helps to sidestep the optimization difficulty of pixel-based representation, which will be discussed later. We use a randomly initialized convolutional U-Net with skip connections for $\mathcal{T}_{\theta}$ and use the randomly sampled code ${\mathbf{z}}{\sim}\mathcal{N}(0,\mathbf{I})\in\mathbb{R}^{H\times W\times 3}$ as a fixed input, where $H$ and $W$ are the height and width of the target texture maps, respectively, and ${\mathbf{z}}$ is fixed during optimization. With this, we re-parameterize the pixel-wise PBR parameters of the texture maps with the neural convolution kernels of the $\mathcal{T}_{\theta}$, i.e., $[\mathbf{K}^{\text{d}}_{\theta},\mathbf{K}^{\text{rm}}_{\theta},\mathbf{K}^{\text{n}}_{\theta}]=\mathcal{T}_{\theta}({\mathbf{z}})$, where $\mathbf{K}^{\text{d}}_{\theta},\mathbf{K}^{\text{n}}_{\theta}\in\mathbb{R}^{H\times W\times 3}$, $\mathbf{K}^{\text{rm}}_{\theta}\in\mathbb{R}^{H\times W\times 2}$, and thus $\mathcal{T}_{\theta}({\mathbf{z}})\in\mathbb{R}^{H\times W\times(3+2+3)}$. We analyze and discuss the effects of the proposed neural re-parameterization in Sec. 4.
@@ 3.4 Text-driven DC-PBR Optimization
$\theta^{*}{=}\operatornamewithlimits{\arg\min}_{\theta}\mathbb{E}_{t,\epsilon}\left[\lVert\hat{\epsilon}_{\phi}(\mathcal{R}^{\mathbf{M}}_{t}(\mathbf{K}^{\text{d}}_{\theta},\mathbf{K}^{\text{rm}}_{\theta},\mathbf{K}^{\text{n}}_{\theta});y,t){-}\epsilon\rVert^{2}_{2}\right],$ (3)
@@ 4 Analysis: Effect of the Deep Convolutional Re-parameterization for PBR Texture Maps
In this section, we discuss the effects and advantages of our parameterization, DC-PBR. We found that the DC-PBR has a favorable learning behavior in frequency. We describe the experiment setup for the analysis in Sec. 4.1 and extend it to the text-driven texture map synthesis task in Sec. 4.2.
@@ Methods
Figures 4a and 4b show the energy-iteration plot of both methods. While the pixel value optimization fits all the frequency bands simultaneously (Fig. 4a), the neural re-parameterized optimization fits the low-frequency components faster and defers to fit the high-frequency components later (Fig. 4b), i.e., schedules the frequency. Considering that lower-frequency bands mostly contain the content of the image while highest-frequency bands mainly correlate with noises in the image, we hypothesize that the scheduled frequency of neural re-parameterization helps the optimization focus more on the content of the texture map in the earlier iterations. The texture map visualizations show the neural re-parameterized optimization fits the overall texture and skin tones, i.e., low-frequency, first in earlier iterations and details later. A similar observation in the natural image domain is reported in [54, 48], and we further show that the consistent result also holds for the PBR representation. On the other hand, the pixel optimization learns low-to-high frequencies simultaneously, which fits noise and texture signals jointly. This yields undesirable optimization paths that may be harmful for sensitive losses like the SDS loss.
@@ 5 Experiments
In this section, we evaluate Paint-it in several aspects. We first show the high-fidelity qualitative results obtained from Paint-it optimization (Sec. 5.1). We then compare the results with the competing methods, qualitatively and quantitatively (Sec. 5.2). We further discuss the effect of our design choices with ablation studies (Sec. 5.3).
@@
5.1 Qualitative Results
In Fig. 6, we visualize the rendered meshes using Paint-it’s synthesized PBR texture maps for a given text prompt. To show the generalizability of the Paint-it synthesis method, we take the subsets of the large-scale mesh datasets: Objaverse [17] and RenderPeople [3] for general objects and clothed humans. For animals, we obtained the template meshes from the quadruped animal linear mesh model, SMAL [66]. We demonstrate that Paint-it can generate photorealistic and vivid textures with material properties such as a mushroom’s matte surface and a teapot’s metallic surface. Paint-it also synthesizes well-segmented human textures, given an untextured mesh. By leveraging the strong generative prior from the pre-trained text-to-image diffusion model, Paint-it faithfully distinguishes texture parts for skin and cloth. Interestingly, Paint-it can generate pseudo-stereoscopic effects, even though the given mesh surface is flat, e.g., the jewels and gems on a crown. We postulate this effect stems from our DC-PBR, where we synthesize the disentangled material properties, diffuse, roughness & metalness, and perturbing tangent space normals. Paint-it also supports the material control or texture transfer for the same input mesh and the relighting using different environment maps, thanks to the synthesized PBR texture maps (Fig. 2).
@@ 5.2 Comparison with Competing Methods
We also evaluate Paint-it with recent text-driven mesh texture synthesis methods, Latent-Paint [35], Fantasia3D [11], Text2Tex [10], and TEXTure [44]. We synthesize texture maps using each method with the same 3D mesh and text input (see Fig. 6). Paint-it synthesizes more vivid, realistic, and consistent textures, compared to texture inpainting methods, Text2Tex and TEXTure. Specifically, they suffer from texture inconsistencies on the mesh surface and the baked lighting effects. The back-projection of the generated RGB image onto the mesh and the limited diffuse texture representation could be the reason. Latent-Paint synthesizes blurry texture and is also limited in diffuse texture. Fantasia3D disentangles the material properties, similar to ours, but produces inconsistent textures since they optimize per-point texture properties.
@@ 5.3 Ablation studies
We also compare full Paint-it with the method without the proposed DC-PBR. As discussed in Sec. 4, neural re-parameterized optimization naturally embodies the frequency scheduling for synthesizing textures. While baseline optimization (Eq. (4)) adopts a regularization term, $\mathcal{L}_{\text{TV}}$, to avoid synthesizing noisy textures with high frequencies, it still introduces severely jittered textures (see Fig. 7a, w/o Re-param., & Fig. 7c). We believe that our novel DC-PBR optimization can be applied not only to text-driven texture map synthesis tasks but also to other texture map reconstruction [32, 60] or generation tasks [8, 52], too.
@@ Acknowledgment
We thank the members of AMILab [1] and RVH group [2] for their helpful discussions and proofreading. The project was made possible by funding from the Carl Zeiss Foundation. This work is funded by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) - 409792180 (Emmy Noether Programme, project: Real Virtual Humans), and the German Federal Ministry of Education and Research (BMBF): Tübingen AI Center, FKZ: 01IS18039A. Gerard Pons-Moll is a Professor at the University of Tübingen endowed by the Carl Zeiss Foundation, at the Department of Computer Science and a member of the Machine Learning Cluster of Excellence, EXC number 2064/1 – Project number 390727645. K. Youwang and T.-H. Oh were supported by Institute of Information & communications Technology Planning & Evaluation (IITP) grant funded by the Korea government(MSIT) (No.RS-2023-00225630, Development of Artificial Intelligence for Text-based 3D Movie Generation; No.RS-2022-00164860, Development of human digital twin technology based on dynamic behavior modeling and human-object-space interaction; No.2021-0-02068, Artificial Intelligence Innovation Hub).
@@ A.3 Details of Frequency-based Analysis
A.4 Details of User Study
We conducted a user study to assess the realism of the synthesized PBR texture maps. We showed ten untextured meshes from the Objaverse [17] dataset and showed five different results obtained from the methods: Latent-Paint [35], Fantasia3D [11], Text2Tex [10], TEXTure [44], and Paint-it (ours). We asked 30 users to rate the realism of the rendered images in the score range 1 to 5, i.e., 1: very unrealistic, 2: unrealistic, 3: neutral, 4: realistic, and 5: very realistic. The order of the methods was randomly shuffled for fairness. The interface of the user study is shown in Fig. S3.
A.5 Optimization Time
The optimization time for synthesizing PBR texture maps takes about 15 min. for general object meshes in Objaverse [17]. For more complex cases, such as 3D humans or animals, we additionally use face-focused mesh renderings; thus, it takes about 30 min. to complete. We use a single NVIDIA RTX A6000 GPU for the optimization.
@@ B Additional Experiment
Similarly, for our task, i.e., text-driven DC-PBR optimization, U-Net w/o skip connection fails to increase the mid-to-high frequency band (purple line), resulting in blurry texture maps. In contrast, our DC-PBR, parameterized in U-Net+skip, successfully increases the mid-to-high frequency band, synthesizing fine-grained texture maps. We conclude that the skip connections are in charge of synthesizing fine-grained, high-frequency details of texture maps. This observation aligns with the Deep Image Prior’s claim, where skip connections benefit the inverse problems with multi-scale feature awareness. We additionally showed the frequency level behavior of the skip connection through the experiments.
@@ C Additional Results
C.1 More Qualitative Results
@@
C.2 More Comparison Results
In Figs. S8 and S9, we provide more qualitative results that compare Paint-it and recent competing methods [35, 11, 10, 44]. As in Fig. 6 of the main paper, we synthesize texture maps using each method for the same untextured meshes and text prompts. Overall, Paint-it synthesizes much realistic and vivid texture on the meshes, thanks to the PBR texture representation and texture smoothness induced by our proposed DC-PBR. Note that Fantasia3D [11] also synthesizes PBR materials, but in a per-point independent manner; thus, it lacks texture smoothness and yields substantial jitterings. Moreover, given an untextured mesh, Fantasia3D converts the mesh into a signed distance function (SDF) representation, DMTet [47]. Such auxiliary re-meshing optimization leads to severe geometric quality degradation, e.g., floating artifacts on a toy bicycle example, in Fig. S8.
C.3 Paint-it for Animated Meshes

### v2

@@ Abstract
11footnotetext: Work done during a visiting research period at the University of Tübingen.
@@ 1 Introduction
To reduce such burdens, methods for generating diverse 3D assets have been extensively studied [40, 58, 11, 34, 41, 8, 17]. Notably, the recent progress in neural volumetric representations, e.g., NeRF [38] and diffusion models [46, 47] have advanced the development of text-driven 3D asset generation [28, 42, 26, 33], which leverages a cheaper guidance, i.e., text description. While these methods generate coarse geometries and textures, the qualities are still unsatisfactory. Moreover, to use the generated assets in real graphics engines, e.g., Blender, one must convert the implicit volumetric geometries and textures into the explicit mesh and compatible texture maps, which makes them impractical. Manual extraction of mesh surfaces and unwrapping of textures could be performed, but it still has limitations. The unwrapped texture maps inevitably have heterogeneous texture mappings, so we cannot easily transfer or edit them, which is crucial for generating diverse 3D assets.
@@
In experiments, we demonstrate that Paint-it produces high-quality texture maps for general 3D meshes, e.g., objects, humans, and animals (Fig. 1). Also, Paint-it synthesizes a remarkable quality of texture maps compared to competing methods. As a favorable byproduct, the synthesized PBR texture maps are compatible with the popular graphics engine and can be seamlessly integrated into relighting and material control pipelines (Fig. 2). We summarize our main contributions as follows.
@@ 3 Paint-it: Text-Driven PBR Texture Synthesis via Neural Re-parameterized Optimization
Figure 3: Paint-it: Overall pipeline. Given a 3D object mesh without a texture and a text describing the desired appearance of the mesh, Paint-it synthesizes realistic PBR texture maps via synthesis-through-optimization. We introduce DC-PBR, which parameterizes the PBR texture map into randomly initialized U-Net convolutional neural kernels. By performing texture mapping to texturize the given mesh, we differentiably rasterize the textured mesh and obtain multi-view images, then compute the diffusion-guided loss. Note that Paint-it optimizes the neural parameters of the U-Net rather than directly optimizing the pixel values of the texture map.
@@ 3.3 DC-PBR: Deep Convolutional PBR Texture Map Re-parameterization
We propose the deep convolutional re-parameterization of PBR texture maps, DC-PBR, $\mathcal{T}_{\theta}$. Instead of the pixel value parameterization of texture maps, using DC-PBR helps to sidestep the optimization difficulty of pixel-based representation, which will be discussed later in Sec. 4. We use a randomly initialized convolutional U-Net with skip connections for $\mathcal{T}_{\theta}$ and use the randomly sampled code ${\mathbf{z}}{\sim}\mathcal{N}(0,\mathbf{I})\in\mathbb{R}^{H\times W\times 3}$ as a fixed input, where $H$ and $W$ are the height and width of the target texture maps, respectively, and ${\mathbf{z}}$ is fixed during optimization. With this, we re-parameterize the pixel-wise PBR parameters of the texture maps with the neural convolution kernels of the $\mathcal{T}_{\theta}$, i.e., $[\mathbf{K}^{\text{d}}_{\theta},\mathbf{K}^{\text{rm}}_{\theta},\mathbf{K}^{\text{n}}_{\theta}]=\mathcal{T}_{\theta}({\mathbf{z}})$, where $\mathbf{K}^{\text{d}}_{\theta},\mathbf{K}^{\text{n}}_{\theta}\in\mathbb{R}^{H\times W\times 3}$, $\mathbf{K}^{\text{rm}}_{\theta}\in\mathbb{R}^{H\times W\times 2}$, and thus $\mathcal{T}_{\theta}({\mathbf{z}})\in\mathbb{R}^{H\times W\times(3+2+3)}$.
@@ 3.4 Text-driven DC-PBR Optimization
$\small\theta^{*}{=}\operatornamewithlimits{\arg\min}_{\theta}\mathbb{E}_{t,\epsilon}\left[\lVert\hat{\epsilon}_{\phi}(\mathcal{R}^{\mathbf{M}}_{t}(\mathbf{K}^{\text{d}}_{\theta},\mathbf{K}^{\text{rm}}_{\theta},\mathbf{K}^{\text{n}}_{\theta});y,t){-}\epsilon\rVert^{2}_{2}\right],$ (3)
@@ Methods
Figures 4a and 4b show the energy-iteration plot of both methods. While the pixel value optimization fits all the frequency bands simultaneously (Fig. 4a), the neural re-parameterized optimization fits the low-frequency components faster and defers to fit the high-frequency components later (Fig. 4b), i.e., schedules the frequency. Considering that lower-frequency bands mostly contain the content of the image while highest-frequency bands mainly correlate with noises in the image, we hypothesize that the scheduled frequency of neural re-parameterization helps the optimization focus more on the content of the texture map in the earlier iterations. The texture map visualizations show the neural re-parameterized optimization fits the overall texture and skin tones, i.e., low-frequency, first in earlier iterations and details later. A similar observation in the natural image domain is reported in [54, 49], and we further show that the consistent result also holds for the PBR representation. On the other hand, the pixel optimization learns low-to-high frequencies simultaneously, which fits noise and texture signals jointly. This yields undesirable optimization paths that may be harmful for sensitive losses like the SDS loss, which is further investigated as follows.
@@ 5 Experiments
5.1 Qualitative Results
In Fig. 6, we visualize the rendered meshes using Paint-it’s synthesized PBR texture maps for a given text prompt. To show the generalizability of the Paint-it synthesis method, we take the subsets of the large-scale mesh datasets: Objaverse [16] and RenderPeople [3] for general objects and clothed humans. For animals, we obtained the template meshes from the quadruped animal linear mesh model [65]. Paint-it can generate photorealistic and vivid textures with material properties such as a mushroom’s matte surface and a teapot’s metallic surface. By leveraging the strong generative prior from the pre-trained text-to-image diffusion model, Paint-it faithfully distinguishes texture parts for skin and cloth. Interestingly, Paint-it can generate pseudo-stereoscopic effects, even though the given mesh surface is flat, e.g., the jewels and gems on a crown. We postulate this effect stems from our DC-PBR, where we synthesize the disentangled material properties along with perturbing tangent space normals. Paint-it also supports the material control or texture transfer for the same input mesh and the relighting using different environment maps, thanks to the synthesized PBR texture maps (Fig. 2).
@@ 5.2 Comparison with Competing Methods
In Fig. 6, we evaluate Paint-it with recent text-driven mesh texture synthesis methods, Latent-Paint [35], Fantasia3D [10], Text2Tex [9], and TEXTure [45]. Texture maps are generated from each method on identical 3D meshes and text. Paint-it synthesizes more vivid, realistic, and consistent textures, compared to texture inpainting methods, Text2Tex and TEXTure. Specifically, they suffer from multi-view texture inconsistencies and the baked lighting effects. The back-projection of the generated RGB image onto the mesh and the limited diffuse texture representation could be the reason. Latent-Paint synthesizes blurry texture and is also limited in diffuse texture. Fantasia3D learns a coordinate-based MLP to predict the per-point PBR materials, whereas we parameterize the full texture map globally. When backpropagating gradients, Paint-it has a global effect over the full texture, while Fantasia3D is much more local. Given that SDS is an ambiguous and noisy signal, the global gradient update of Paint-it helps get a higher-quality, coherent appearance. In Fig. 7, Paint-it obtain better-disentangled materials and specular properties, while Fantasia3D fails to generate the mug’s metallic (smooth) surface. Also, Fantasia3D re-meshes the input mesh, destroying the geometry and obtaining implausible uv mapping with substantial seams.
@@ 5.3 Ablation studies
As discussed in Sec. 4, DC-PBR, i.e., neural re-parameterized optimization, naturally embodies the frequency scheduling for synthesizing textures. While baseline optimization (Eq. (4)) adopts a regularization term to avoid synthesizing noisy textures with high frequencies, it still introduces severely jittered textures (see Fig. 8a, w/o Re-param., & Fig. 8c).
@@ Acknowledgment
We thank the members of AMILab [1] and RVH group [2] for their helpful discussions and proofreading. The project was made possible by funding from the Carl Zeiss Foundation. This work is funded by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) - 409792180 (Emmy Noether Programme, project: Real Virtual Humans), and the German Federal Ministry of Education and Research (BMBF): Tübingen AI Center, FKZ: 01IS18039A. Gerard Pons-Moll is a Professor at the University of Tübingen endowed by the Carl Zeiss Foundation, at the Department of Computer Science and a member of the Machine Learning Cluster of Excellence, EXC number 2064/1 – Project number 390727645. Kim Youwang and Tae-Hyun Oh were supported by Institute of Information & communications Technology Planning & Evaluation (IITP) grant funded by the Korea government(MSIT) (No.RS-2023-00225630, Development of Artificial Intelligence for Text-based 3D Movie Generation; No.2022-0-00290, Visual Intelligence for Space-Time Understanding and Generation based on Multi-layered Visual Common Sense; No.2021-0-02068, Artificial Intelligence Innovation Hub).
@@ A.3 Details of Frequency-based Analysis
In Fig. 4-left, the optimization starts from a monotonic gray image, which has high zero-frequency energies, i.e., DC component. However, note that we visualize each frequency band’s ‘average’ energy. DC component occupies only a tiny area ($\ll 1\%$) in the lowest frequency band; thus, the pixel brightness hardly affects the band’s energy.
@@
Figure S4: Feature grid+MLP (baseline) vs. DC-PBR (ours)
@@
A.4 Details of User Study
We conducted a user study to assess the realism of the synthesized PBR texture maps. We showed ten untextured meshes from the Objaverse [16] dataset and showed five different results obtained from the methods: Latent-Paint [35], Fantasia3D [10], Text2Tex [9], TEXTure [45], and Paint-it (ours). We asked 30 users with engineering/non-engineering backgrounds to rate the realism of the rendered images in the score range 1 to 5, i.e., 1: very unrealistic, 2: unrealistic, 3: neutral, 4: realistic, and 5: very realistic. The order of the methods was randomly shuffled for fairness. The interface of the user study is shown in Fig. S3.
A.5 Details & Discussion about Optimization Time
The optimization time for synthesizing PBR texture maps takes about 15 min. for general object meshes in Objaverse [16]. For more complex cases, such as 3D humans or animals, we additionally use face-focused mesh renderings; thus, it takes about 30 min. to complete. We optimized both the baseline (Eq. (4) in the main paper, pixel optimization) and our DC-PBR until convergence, and there’s no significant time difference. Instead, DC-PBR helps the optimization with noisy SDS loss find a better solution than the vanilla pixel optimization. We use a single NVIDIA RTX A6000 GPU for the optimization.
Extending Paint-it to large-scale 3D scenes would be an interesting future direction. However, Paint-it’s PBR texture generation for large-scale scenes would take a longer optimization time. As suggested in FPRF [29], we may try semantic feature distillation to accelerate the optimization when stylizing large-scale 3D scenes using SDS.
@@ B Additional Experiment
B.1 More Baseline for PBR Representation
For a PBR representation baseline, other than direct pixel optimization (Eq. (4) in the main paper), we implement a multi-resolution hash encoding of grid features and subsequent MLP [39] to model the disentangled PBR texture maps. In Fig. S4, our DC-PBR representation yields smoother and more vivid texture results than the new baseline. In Table S1, our DC-PBR obtains a better FID score than other PBR representation baselines (Pixel optim., & Feat. grid+MLP). We postulate that the SDS gradients only propagate to local hash grids in the new baseline, lacking non-local texture smoothness, as supported in [32]. On the other hand, the spatial CNN kernels of our DC-PBR are beneficial in naturally imposing texture smoothness.
B.2 Effects of DC-PBR Design Choice
@@
Similarly, for our task, i.e., text-driven DC-PBR optimization, U-Net w/o skip connection fails to increase the mid-to-high frequency band (purple line), resulting in blurry texture maps. In contrast, our DC-PBR, parameterized in U-Net+skip, successfully increases the mid-to-high frequency band, synthesizing fine-grained texture maps. We conclude that the skip connections are in charge of synthesizing fine-grained, high-frequency details of texture maps. This observation aligns with the Deep Image Prior’s claim, where skip connections benefit the inverse problems with multi-scale feature awareness. We additionally showed the frequency level behavior of the skip connection through the experiments (see Fig. S6).
Pixel optim. Only $\mathbf{K}^{\text{d}}$ Feat. grid+MLP DC-PBR (Ours)
FID ($\downarrow$) 216.6 59.39 55.98 34.46
Table S1: FID for ablation study
@@ C Additional Results
C.1 More Quantitative Comparison
In Tab. S1, we report FID scores for the ablation study (in Sec. 5.3 in the main paper). We used 410 meshes from Objaverse to get the real sample set and added 50 meshes for the generated sample set. The results support that our full DC-PBR enhances the realism of the generated texture.
C.2 More Qualitative Results
@@
C.3 More Comparison Results
In Figs. S10 and S11, we provide more qualitative results that compare Paint-it and recent competing methods [35, 10, 9, 45]. As in Fig. 6 of the main paper, we synthesize texture maps using each method for the same untextured meshes and text prompts. Overall, Paint-it synthesizes much realistic and vivid texture on the meshes, thanks to the PBR texture representation and texture smoothness induced by our proposed DC-PBR. Note that Fantasia3D [10] also synthesizes PBR materials, but in a per-point independent manner; thus, it lacks texture smoothness and yields substantial jitterings. Moreover, given an untextured mesh, Fantasia3D converts the mesh into a signed distance function (SDF) representation, DMTet [48]. Such auxiliary re-meshing optimization leads to severe geometric quality degradation, e.g., floating artifacts on a toy bicycle example, in Fig. S10.
C.4 Paint-it for Animated Meshes